OpenStudy (anonymous):

If \(\color{green}{pth \; , \; \color{blue}{qth}}\) and \(\color{red}{rth}\) terms of a \(\color{green}{\mathbf{G} \text{eometric} \; \mathbf{P} \text{rogression}}\) be : \(\color{blue}{a, b}\) and \(\color{blue}{c}\) respectively, then: the value of : \[\color{green}{\large (a^{q -r} \times b^{r-p} \times c^{p-q}) } \; \; is :\] \(\color{red}{Answer \; : \; 1}\)

OpenStudy (anonymous):

is it \[\frac{ab}{c^2}\]

hartnn (hartnn):

yes,i am getting 1 :)

OpenStudy (anonymous):

Please explain @hartnn

OpenStudy (anonymous):

lol...its \[1\]

hartnn (hartnn):

let F be 1st term,R be common ratio....a=F*R^(p-1),similarly,b,c....question asked can be expressed as (a/c)^q*(b/a)^r*(c/a)^p.....(a/c)^q=R^(pq-rq)......adding all powers of R we get 0,so ans=1......sorry i m not good at explaining neatly with cool pics :P

OpenStudy (anonymous):

I show you clearly what @hartnn meant. \[\large pth \; term = A \cdot r^{p-1} \implies a = A \cdot r^{p-1}\] \[\large qth \; term = A \cdot r^{q-1} \implies b = A \cdot r^{q-1}\] \[\large rth \; term = A \cdot r^{r-1} \implies a = A \cdot r^{r-1}\] \[a^{q -r} \times b^{r-p} \times c^{p-q} \implies A^{(q-r +r-p+p-q)} \cdot r^{(p-1)(q-r)+(q-1)(r-p)+(r-1)(p-q)}\] \[\implies A^{0} \cdot r^{(pq -pr -q + r +rq -pq -r + p + pr -rq -p + q)} \implies A^{0} \cdot r^{0}\] \[\implies (1) \cdot(1) = 1\]

hartnn (hartnn):

how to write like that?...not using draw......using equation??how?

OpenStudy (anonymous):

Using equation button..

hartnn (hartnn):

i was trying to write upon sign...like a/c...how do i make a and c show in different line with dash in between?

OpenStudy (anonymous):

Like: \[\frac{a}{c}\]

OpenStudy (anonymous):

??

hartnn (hartnn):

yes,how u did that?

OpenStudy (anonymous):

\frac{}{}

OpenStudy (anonymous):

In equation editor write; frac{}{} brackets must be these only.. numerator will in first bracket and denominator will be in second one.. frac{a}{c}

hartnn (hartnn):

i have to write { then f then r then a....no shortcut??

OpenStudy (anonymous):

I mean it is a code..

hartnn (hartnn):

\[\frac{a}{c}\]

hartnn (hartnn):

hmmmm....tedious :P

OpenStudy (anonymous):

Great.. Can you write this : a/b/c/d ??

hartnn (hartnn):

\[\frac{a}{b}{c}{d}\]

OpenStudy (anonymous):

I wish I could give you more medals for this.. Ha ha ha..

hartnn (hartnn):

frac{a}{b}{c}{d} doesn't work....

OpenStudy (anonymous):

It won't You can use {}{} two times only not 4 times..

hartnn (hartnn):

\[\frac{\frac{a}{b}}{\frac{c}{d}}\]

hartnn (hartnn):

frac{frac{}{}{frac{}{}} worked :)

OpenStudy (anonymous):

Yes it will.. Great work.. Now one I thing I must do for you..

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To The World Of Latex..}}\]

hartnn (hartnn):

oh,so this is latex!heard about it.....

OpenStudy (anonymous):

@hartnn If you want to learn LATEX then you can post your question here : http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)

hartnn (hartnn):

ohh!!separate latex group......'LATEXIANS'... :P

OpenStudy (anonymous):

Yes..